Advertisements
Advertisements
प्रश्न
Suppose S is the event that will happen tomorrow and P(S) = 0.03. Find P(S’)
Advertisements
उत्तर
Given, P(S) = 0.03
P(S’) = 1 – P(S)
P(S’) = 1 – 0.03 ............[∵ P(S) + P(S’) = 1]
P(S’) = 0.97
APPEARS IN
संबंधित प्रश्न
A coin is tossed 80 times and the head is obtained 38 times. Now, if a coin tossed once, what will the probability of getting a tail
A coin is tossed 80 times and the head is obtained 38 times. Now, if a coin tossed once, what will the probability of getting a head
There are 8 marbles in a bag with numbers from 1 to 8 marked on each of them. What is the probability of drawing a marble with a number
(i) 3
(ii) 7
Two coins are tossed simultaneously 100 times and the outcomes are as given below:
| Outcomes | Two heads (H, H) |
Exactly one head (H T or T H) |
No head (T T) |
| No. of times | 21 | 55 | 24 |
If the same pair of coins is tossed again at random, find the probability of getting:
(i) two heads
(ii) exactly one head
(iii) no head.
A bag contains 4 white and 6 black balls,- all of the same shape and same size. A ball is drawn from the bag without looking into the bag. Find the probability that the ball drawn is:
(i) a black ball
(ii) a white ball
(iii) not a black ball
In a single throw of a dice, find the probability of getting a number 4
Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 50
Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 80
Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 40
A carton contains eight brown and four white eggs. Find the probability that an egg selected at random is:
(i) brown
(ii) white
